Solve (x + 5)^2 = 3
Can anyone help me on this please? Thank you so much!

Answer by hailu(13)

(x+5)(x+5)=3
x²+5x+5x+25=3
x²+10x+25=3
=> get the variables,and no.s on opposite sides:
x²+10x=3-25
x²+10x= -22
=> then change it into a quadratic equation,add +22 on both sides to equal zero
x²+10x+22=0
=>take the -22 back where it was, to expand on one side, and add on the other
x²+5x+5x+25= -22+25
=>factor each one
x(x+5)+5(x+5)= 3
(x+5)²=3
(x+5)=±√(3)
=> your answer will be
x= -5±√(3)
=> which gives u two answers, either:
x=-5+√(3)
=> OR
x=-5-√(3)
hope u got it!
name's HaNaWi, owe my comprehention 2 ma best teacher, hailu